Landon can climb a certain hill at a rate that is 3.5 mph slower than his rate coming down the hill. If it take him 4 hours to climb the hill and 135 minutes to come down the hill. What is his rate coming down?

Answer by richwmiller(17219)

x*136/60=d,
(x-3.5)*3=d
x*136/60=(x-3.5)*3
x*136/60=3x-10.5
x*136/6=30x-105
x*136=180x-630
630=44x
315=22x
315/22=x
x=14.32 mph downhill
x-3.5=10.82 mph uphill